# Limb Observations of Global Solar Coronal EUV Wavefronts: the Inclination, Kinematics, Coupling with the Expanding CMEs, and Connection with the CME-driven Shocks

### v1

@@ Abstract
We select and investigate six global solar extreme ultraviolet (EUV) wave events using data from the Solar Dynamics Observatory (SDO) and the Solar and Heliospheric Observatory (SOHO). These eruptions are all on the limb but recorded as halo coronal mass ejections (CMEs) because the CME-driven shocks have expanded laterally to the opposite side. With the limb observations avoiding the projection effect, we have measured the inclination and speed of the EUV wavefront from 1.05 to 1.25 . We also investigate the coupling and connection of the EUV wavefront with the CME boundary and the CME-driven shock, respectively. The major findings in the six events are: (1) the forward inclination of the primary and coronal-hole transmitted EUV wavefronts is estimated, respectively, and the origins of these inclinations and their effects on the estimate of actual wavefront speed are investigated; (2) the wavefront speed can be elevated by loop systems near the coronal base, and the average speed in the low corona has no clear correlation with the lateral expansion of the CME-driven shock in the high corona; (3) the fast magnetosonic Mach number of the wavefront is larger than unity from the coronal base; (4) the EUV wavefront is coupled with the CME driver throughout the propagation in two events; (5) after the EUV wavefront vanishes, the CME-driven shock continues traveling on the opposite side and disconnects from the EUV wavefront in four events. These results and their implications are discussed, which provide insight into the properties of global EUV waves.
@@ 1 Introduction
A CME-driven shock observed in white light in the outer corona is often connected with an EUV wave, as a footprint of the shock in the low corona (e.g., Veronig et al., 2010; Cheng et al., 2012; Kwon et al., 2014; Zhu et al., 2018; Liu et al., 2019c). This connection indicates the shock nature of the corresponding EUV wave and the global geometry of the shock. Most solar type II radio bursts are signatures of CME-driven shocks (e.g., Uchida, 1960; Mann et al., 1995; Claßen & Aurass, 2002; Liu et al., 2009; Cho et al., 2011; Kumari et al., 2023), and some are attributed to non-CME generated coronal shocks (e.g., Magdalenić et al., 2010; Su et al., 2015; Hou et al., 2023). For a CME-driven shock, a type II radio burst can be used to infer the shock’s propagation properties (e.g., Reiner et al., 2007; Liu et al., 2008, 2013; Cremades et al., 2015; Hu et al., 2016; Zhao et al., 2019). A CME-driven shock can propagate laterally to the other side of the Sun opposite the eruption site, which has been observed with white light in the outer corona (Kwon et al., 2014; Liu et al., 2017; Hu et al., 2019). Henceforth, for the sake of brevity, we refer to a “CME-driven shock observed in white light” as a “white-light shock”. In the contribution of the transmission of two polar coronal holes, the EUV wave on 2017 September 10 travels all over the solar sphere and reaches the opposite side of the eruption (e.g., Liu et al., 2018; Hu et al., 2019), which is unique. Although the eruption is on the west limb, the CME manifests as a halo because the bubble of the white-light shock encloses the Sun (similar to cases in Kwon et al., 2014). The wavefront of this EUV wave, throughout the propagation, is continuously spatially connected with the white-light shock even on the opposite side. However, in common events where the CME-driven shock propagates to the opposite side without the transmission of large-scale coronal holes, the connection between the EUV wave and the white-light shock on the opposite side is still elusive.
@@
In this work, we will select and investigate six global EUV wave events which are all triggered by eruptions on the solar limb. The associated CMEs are all recorded as halos because their white-light shocks have propagated to the other side of the Sun opposite the eruption site. This work aims to comprehensively study (1) the local morphology and (2) kinematics of global EUV waves in the low corona, as well as (3) the coupling with the expanding CME and (4) connection with the CME-driven shocks. Observations of limb events can avoid the projection effect, which makes it simpler to obtain the inclination and kinematics of the EUV wavefront. Limb observations are also effective to determine the CME boundary, the EUV wavefront, and the CME-driven white-light shock, which is helpful to discern the decoupling of the EUV wavefront from the CME, as well as the connection between the EUV wavefront and the white-light shock. The data, event selection, and the overview of the events are presented in Section 2. The analysis results are provided in Section 3 and are discussed in Section 4. The conclusions are remarked in Section 5.
@@ 2 Data and Events
All six selected EUV events are associated with halo CMEs recorded in the SOHO/LASCO CME catalog (Gopalswamy et al., 2009b), which are all initiated by eruptions on the solar limb from the view of the Earth. Because high-cadence data are better to acquire the kinematics of EUV waves, especially for those fast ones (Long et al., 2008; Patsourakos & Vourlidas, 2012; Nitta et al., 2013; Downs et al., 2021), we use data of the Atmospheric Imaging Assembly (AIA, Lemen et al., 2012) on board the Solar Dynamics Observatory (SDO, Pesnell et al., 2012) with a cadence of 12 seconds. Although the EUV wave is prominent in running-difference images of AIA 211, 193, and 171 Å, the 211 Å images are chosen because they have a relatively better feature of the EUV wavefront in our events. Below are the selection criteria for the events to be studied: (1) Only halo CMEs with a total width of 360° noted in the CME catalog are selected, which means that each CME-driven shock has propagated to the other side opposite the eruption site; (2) Events with eruption sites at heliographic longitudes between 80° and 100° are considered to be limb events in this study; (3) On the limb an EUV wave has southward and northward wavefronts, of which the wavefront with a larger propagation distance is chosen for investigation in each event; (4) If both the southward and northward wavefronts are too diffuse to be identified or have a propagation distance less than 60° from the eruption site in the position angle, the corresponding events are also excluded. With these criteria, six events have been selected from dozens of events from 2011 September to 2020 November, and are listed in Table 1. Most of these limb events are remarkable and have been studied in prolific literatures. Below in this section we will give an overview for each event, and the detailed results will be presented in the following sections.
@@
The 2017 September 10 EUV wave event: This is a star textbook event, which has been studied in numerous publications. The eruption is at S09W92 in AR 12673 on the west limb, associated with an X8.2 flare peaked at 16:06 UT on 2017 September 10. The flare has been investigated by ground- and space-based instruments at multiple wavelengths (e.g., Seaton & Darnel, 2018; Omodei et al., 2018; Gary et al., 2018; Hayes et al., 2019; Zhao et al., 2021). Turbulent magnetic reconnections are observed in an extended current sheet, which are studied in a plethora of observations or models (e.g., Yan et al., 2018; Li et al., 2018; Cheng et al., 2018; French et al., 2019; Chen et al., 2020; Yu et al., 2020). The CME and its driven shock are both of spectacular speed and scale (e.g., Gopalswamy et al., 2018; Veronig et al., 2018; Hu et al., 2019; Liu et al., 2019c; Yang et al., 2021). The CME-driven shock is also related to type II radio bursts and contributes to the acceleration of energetic particles (e.g., Zhao et al., 2018; Gopalswamy et al., 2018; Morosan et al., 2019; Zhu et al., 2021; de Koning et al., 2022). A series of interplanetary sequences of this event have been observed (e.g., Lee et al., 2018; Cramer et al., 2020; Ding et al., 2020; Kocharov et al., 2020). The EUV wave is currently the only known one that travels all over the solar sphere in the contribution of the transmission of two polar coronal holes (Liu et al., 2018; Hu et al., 2019). In this study, we will investigate the local geometry and kinematics of the EUV wavefront in the low corona, as well as the connection with the associated CME and white-light shock. The southward wavefront is analyzed because it is more prominent than the northward wavefront in AIA 211 Å images.
@@ 3.1 Inclination of the Wavefront
The local morphology of the EUV wavefronts in the low corona on the solar limb is clearly shown in Figure 1. Obvious inclination of the primary wavefronts is observed for five of the six events, except for the 2013 May 13 event. Figure 1 (a)–(e) illustrate a typical frame of the EUV wavefront for the five events, respectively. The dashed-circles mark the altitudes from 1.05 to 1.25 with a step of 0.05 , where time-distance profiles in position angles (measured counterclockwise from the solar North) are extracted. The time-distance profiles for each event are shown in panel (a) of Figures 2–7. If a wavefront is forward inclined, the wavefront at a larger altitude is ahead of that at a smaller altitude, i.e., has a larger distance in the time-distance profile. The time period during which the inclination is observed is indicated by the two vertical dotted lines in panel (a) of Figures 2–7. For each time in the period, the EUV wavefront is determined by a linear fit of the points at the five altitudes from 1.05 to 1.25 . For the 2013 November 7 and 2020 November 29 events, the point at 1.25 is excluded from the fit because the position angle at 1.25 is very close to that at 1.20 (see panel (a) of Figures 4 and 7) indicating the wavefront above 1.20 is not notably inclined. In this study, the inclination angle is defined as the acute angle between the fitted wavefront and the tangent at the average position angle for each time. The average inclination angle is calculated over the time period indicated in panel (a) of Figures 2–7 (except Figure 3 because wavefront inclination is not obvious for the 2013 May 13 event). The average inclination angles for the five events are given in Column (5) of Table 1 and are denoted in Figure 1(a)–(e).
@@ 3.2 Kinematics of the Wavefront
For the 2015 May 5 event, the speed profiles are shown in Figure 5, showing a notable deceleration in the initial stage. Near position angle 10° close to the north pole, the EUV wavefront encounters a coronal hole and the wavefront at lower altitudes is transmitted by the coronal hole. However, the fast magnetosonic speed inside the coronal hole near the north pole (for position angle smaller than 10°) derived from the MHD model is not as high as expected for a coronal hole. This could be due to the large measurement uncertainty of magnetic fields in the polar region. A rapid deceleration of the wavefront at altitudes 1.20 and 1.25 is observed, which also indicates the existence of the coronal hole. As shown in Figure 5(e)–(f), the deceleration occurs a little earlier before the wavefront reaches position angle 10°. It is possibly because the magnetic fields of a coronal hole are diverging, which induces a larger angular width of the coronal hole at higher altitudes (Altschuler et al., 1972; Kopp & Holzer, 1976). The transmitted wavefronts in the coronal hole are discernible only at 1.05 and 1.10 , which has a significantly increased speed. The inclination angle of ° is much smaller than that of the primary wavefront (°, see Column (5) of Table 1). This means that the actual speed of the wavefront in the coronal hole can be lower than the apparent speed. As indicated by the cyan curves in Figure 5(b)–(c), the actual speed calculated by taking into account the inclination is comparable to the speed of the primary wavefront before entering the coronal hole. The primary wavefront is also forward inclined, and the actual speed can also be lower than the apparent speed. Before entering the coronal hole, the Mach number is mostly above 1.5 at all altitudes. However, the Mach number inside the coronal hole is not given in Figure 5 because of the inconsistent fast magnetosonic speed from the MAS MHD model.
@@
For the notable global 2017 September 10 event, the transmission of the EUV wavefront by two polar coronal holes has been reported in previous researches (e.g., Liu et al., 2018; Hu et al., 2019). Before entering the south polar coronal hole, the wavefront speed increases with the altitude from 600 km s-1 to 1000 km s-1 (see Figure 6(b)–(f)). From the inbound boundary of the coronal hole (near position angle 210°), only the wavefront at altitudes of 1.05–1.15 is visible. Inside the coronal hole, the apparent wavefront speed is very high, and ranges from 1000 km s-1 to even 3000 km s-1, which is comparable to the estimate of Liu et al. (2018). After leaving the south polar coronal hole, the wavefront undergoes a gradual deceleration before it collides with the wavefront transmitted by the north polar coronal hole (see Hu et al., 2019). Behind the outbound boundary of the south polar coronal hole (near position angle 140°), the apparent wavefront speed is still km s-1 and the Mach number can be nearly 3 (see Figure 6(b)–(d)). Considering the inclination angle of ° for the transmitted wavefront, the actual speed can be below 500 km s-1 and the Mach number can be below 2 (see the cyan curves in Figure 6(b)–(d)). Inside the coronal hole the inclination angle is not estimated because the wavefront is too diffuse to obtain a reliable angle with visible inspection. If we assume the inclination angle is the same as the transmitted wavefront behind the coronal hole, the actual wavefront speed inside the coronal hole can also be lower than the apparent speed. In the initial stage, the inclination angle of the primary wavefront is °, which means the wavefront is less inclined than those of other events and the actual speed is approximately equal to the apparent speed.
@@ 3.3 Coupling of the CME Boundary and the EUV Wave
For the 2011 September 22 and 2013 November 7 events, we can see the angular distance between the EUV wavefront and the CME boundary is close to zero (see Column (6) of Table 2), which suggests that the two features are coupled until the EUV wave fades out. For the other four events, the angular distance between the EUV wavefront and the CME boundary is larger than 30°, which will be discussed separately below.
@@
Due to the large cadence of the SOHO/LASCO C2 images, it is not possible to track the CME boundary during the whole propagation of the EUV wavefront. For the 2011 September 22 and 2013 November 7 events, the angular distance between the EUV wavefront and the CME boundary is approximately zero, when the wavefront fades out at all altitudes. Although the white-light CME structure below 2 cannot be observed by SOHO/LASCO C2, the CME boundary is almost radial as illustrated in Figure 9 and can extend straight downward to below 2 . This is consistent with that the legs or the overlying loops of the CME boundary are stretched to straight during the rising and expansion of a CME structure. It is indicated that the EUV wavefront is still coupled with the CME boundary when the wavefront vanishes. This suggests during the whole propagation, the EUV wavefront is not propagating freely but is continuously driven by the expanding CME.

### v2

@@ Abstract
We select and investigate six global solar extreme-ultraviolet (EUV) wave events using data from the Solar Dynamics Observatory (SDO) and the Solar and Heliospheric Observatory (SOHO). These eruptions are all on the limb but recorded as halo coronal mass ejections (CMEs) because the CME-driven shocks have expanded laterally to the opposite side. With the limb observations avoiding the projection effect, we have measured the inclination and speed of the EUV wavefront from 1.05 to 1.25 . We also investigate the coupling and connection of the EUV wavefront with the CME boundary and the CME-driven shock, respectively. The major findings in the six events are: (1) the forward inclination of the primary and coronal-hole transmitted EUV wavefronts is estimated, respectively, and the origins of these inclinations and their effects on the estimate of actual wavefront speed are investigated; (2) the wavefront speed can be elevated by loop systems near the coronal base, and the average speed in the low corona has no clear correlation with the lateral expansion of the CME-driven shock in the high corona; (3) the fast magnetosonic Mach number of the wavefront is larger than unity from the coronal base; (4) the EUV wavefront is coupled with the CME driver throughout the propagation in two events; (5) after the EUV wavefront vanishes, the CME-driven shock continues traveling on the opposite side and disconnects from the EUV wavefront in four events. These results and their implications are discussed, which provide insight into the properties of global EUV waves.
@@ 1 Introduction
A CME-driven shock observed in white light in the outer corona is often connected with an EUV wave, the footprint of the shock in the low corona (e.g., Veronig et al., 2010; Cheng et al., 2012; Kwon et al., 2014; Zhu et al., 2018; Liu et al., 2019c). This connection indicates the shock nature of the corresponding EUV wave and the global geometry of the shock. Most solar type II radio bursts are signatures of CME-driven shocks (e.g., Uchida, 1960; Mann et al., 1995; Claßen & Aurass, 2002; Liu et al., 2009; Cho et al., 2011; Kumari et al., 2023), and some are attributed to non-CME generated coronal shocks (e.g., Magdalenić et al., 2010; Su et al., 2015; Hou et al., 2023). For a CME-driven shock, a type II radio burst can be used to infer the shock’s propagation properties (e.g., Reiner et al., 2007; Liu et al., 2008, 2013; Cremades et al., 2015; Hu et al., 2016; Zhao et al., 2019). A CME-driven shock can propagate laterally to the other side of the Sun opposite the eruption site, which has been observed with white light in the outer corona (Kwon et al., 2014; Liu et al., 2017; Hu et al., 2019). Henceforth, for the sake of brevity, we refer to a “CME-driven shock observed in white light” as a “white-light shock”. In the contribution of the transmission of two polar coronal holes, the EUV wave on 2017 September 10 travels all over the solar sphere and reaches the opposite side of the eruption (e.g., Liu et al., 2018; Hu et al., 2019), which is unique. Although the eruption is on the west limb, the CME manifests as a halo because the bubble of the white-light shock encloses the Sun (similar to cases in Kwon et al., 2014). The wavefront of this EUV wave, throughout the propagation, is continuously spatially connected with the white-light shock even on the opposite side. However, in common events where the CME-driven shock propagates to the opposite side without the transmission of large-scale coronal holes, the connection between the EUV wave and the white-light shock on the opposite side is still elusive.
@@
In this work, we will select and investigate six global EUV wave events which are all triggered by eruptions on the solar limb. The associated CMEs are all recorded as halos because their white-light shocks have propagated to the other side of the Sun opposite the eruption site. This work aims to comprehensively study (1) the local morphology and (2) kinematics of global EUV waves in the low corona, as well as (3) the coupling with the expanding CMEs and (4) connection with the CME-driven shocks. Observations of limb events can avoid the projection effect, which makes it simpler to obtain the inclination and kinematics of the EUV wavefront. Limb observations are also effective to determine the CME boundary, the EUV wavefront, and the CME-driven white-light shock, which is helpful to discern the decoupling of the EUV wavefront from the CME, as well as the connection between the EUV wavefront and the white-light shock. The data, event selection, and the overview of the events are presented in Section 2. The analysis results are provided in Section 3 and are discussed in Section 4. The conclusions are remarked in Section 5.
@@ 2 Data and Events
All six selected EUV events are associated with halo CMEs recorded in the SOHO/Large Angle and Spectrometric Coronagraph Experiment (LASCO) CME catalog (Gopalswamy et al., 2009b), which are all initiated by eruptions on the solar limb from the view of the Earth. Because high-cadence data are better to acquire the kinematics of EUV waves, especially for those fast ones (Long et al., 2008; Patsourakos & Vourlidas, 2012; Nitta et al., 2013; Downs et al., 2021), we use data of the Atmospheric Imaging Assembly (AIA, Lemen et al., 2012) on board the Solar Dynamics Observatory (SDO, Pesnell et al., 2012) with a cadence of 12 seconds. Although the EUV wave is prominent in running-difference images of AIA 211, 193, and 171 Å, the 211 Å images are chosen because they have a relatively better feature of the EUV wavefront in our events. Below are the selection criteria for the events to be studied: (1) Only halo CMEs with a total width of 360° noted in the CME catalog are selected, which means that each CME-driven shock has propagated to the other side opposite the eruption site; (2) Events with eruption sites at heliographic longitudes between 80° and 100° are considered to be limb events in this study; (3) On the limb an EUV wave has southward and northward wavefronts, of which the wavefront with a larger propagation distance is chosen for investigation in each event; (4) If both the southward and northward wavefronts are too diffuse to be identified or have a propagation distance less than 60° from the eruption site in the position angle, the corresponding events are also excluded. With these criteria, six events have been selected from dozens of limb events from 2011 September to 2020 November, and are listed in Table 1. Most of these limb events are remarkable and have been studied in prolific literatures. Below in this section we will give an overview for each event, and the detailed results will be presented in the following sections.
@@
The 2017 September 10 EUV wave event: This is a star textbook event, which has been studied in numerous publications. The eruption is at S09W92 in AR 12673 on the west limb, associated with an X8.2 flare peaked at 16:06 UT on 2017 September 10. The flare has been investigated by ground- and space-based instruments at multiple wavelengths (e.g., Seaton & Darnel, 2018; Omodei et al., 2018; Gary et al., 2018; Hayes et al., 2019; Zhao et al., 2021). Turbulent magnetic reconnections are observed in an extended current sheet, which have been studied in a plethora of observations or models (e.g., Yan et al., 2018; Li et al., 2018; Cheng et al., 2018; French et al., 2019; Chen et al., 2020; Yu et al., 2020). The CME and its driven shock are both of spectacular speed and scale (e.g., Gopalswamy et al., 2018; Veronig et al., 2018; Hu et al., 2019; Liu et al., 2019c; Yang et al., 2021). The CME-driven shock is also related to type II radio bursts and contributes to the acceleration of energetic particles (e.g., Zhao et al., 2018; Gopalswamy et al., 2018; Morosan et al., 2019; Zhu et al., 2021; de Koning et al., 2022). A series of interplanetary sequences of this event have been observed (e.g., Lee et al., 2018; Cramer et al., 2020; Ding et al., 2020; Kocharov et al., 2020). The EUV wave is currently the only known one that travels all over the solar sphere in the contribution of the transmission of two polar coronal holes (Liu et al., 2018; Hu et al., 2019). In this study, we will investigate the local geometry and kinematics of the EUV wavefront in the low corona, as well as the connection with the associated CME and white-light shock. The southward wavefront is analyzed because it is more prominent than the northward wavefront in AIA 211 Å images.
@@ 3.1 Inclination of the Wavefront
The local morphology of the EUV wavefronts in the low corona on the solar limb is clearly shown in Figure 1. Obvious inclination of the primary wavefronts is observed for five of the six events, except for the 2013 May 13 event. Figure 1 (a)–(e) illustrate a typical frame of the EUV wavefront for the five events, respectively. The dashed-circles mark the altitudes from 1.05 to 1.25 with a step of 0.05 , where time-distance profiles in position angles (measured counterclockwise from the solar North) are extracted. The time-distance profiles for each event are shown in panel (a) of Figures 2–7. If a wavefront is forward inclined, the wavefront at a larger altitude is ahead of that at a smaller altitude, i.e., has a larger distance in the time-distance profile. The time period during which the inclination is observed is indicated by the two vertical dotted lines in panel (a) of Figures 2–7 (except Figure 3 because wavefront inclination is not obvious for the 2013 May 13 event). For each time in the period, the EUV wavefront is determined by a linear fit of the points at the five altitudes from 1.05 to 1.25 . For the 2013 November 7 and 2020 November 29 events, the point at 1.25 is excluded from the fit because the position angle at 1.25 is very close to that at 1.20 (see panel (a) of Figures 4 and 7) indicating the wavefront above 1.20 is not notably inclined. In this study, the inclination angle is defined as the acute angle between the fitted wavefront and the tangent at the average position angle for each time. The average inclination angle is calculated over the time period indicated in panel (a) of Figures 2–7 (except Figure 3). The average inclination angles for the five events are given in Column (5) of Table 1 and are denoted in Figure 1(a)–(e).
@@ 3.2 Kinematics of the Wavefront
For the 2015 May 5 event, the speed profiles are shown in Figure 5, showing a notable deceleration in the initial stage. Near position angle 10° close to the north pole, the EUV wavefront encounters a coronal hole and the EUV wave is transmitted by the coronal hole. However, the fast magnetosonic speed inside the coronal hole near the north pole (for position angles smaller than 10°) derived from the MHD model is not as high as expected for a coronal hole. This could be due to the large measurement uncertainty of magnetic fields in the polar region. A rapid deceleration of the wavefront at altitudes 1.20 and 1.25 is observed, which also indicates the existence of the coronal hole. As shown in Figure 5(e)–(f), the deceleration occurs a little earlier before the wavefront reaches position angle 10°. It is possibly because the magnetic fields of a coronal hole are diverging, which induces a larger angular width of the coronal hole at higher altitudes (Altschuler et al., 1972; Kopp & Holzer, 1976). The transmitted wavefront in the coronal hole is discernible only at 1.05 and 1.10 , which has a significantly increased speed. The inclination angle of ° is much smaller than that of the primary wavefront (°, see Column (5) of Table 1). This means that the actual speed of the wavefront in the coronal hole can be lower than the apparent speed. As indicated by the cyan curves in Figure 5(b)–(c), the actual speed calculated by taking into account the inclination is comparable to the speed of the primary wavefront before entering the coronal hole. The primary wavefront is also forward inclined, and the actual speed can also be lower than the apparent speed. Before entering the coronal hole, the Mach number is mostly above 1.5 at all altitudes. However, the Mach number inside the coronal hole is not given in Figure 5 because of the inconsistent fast magnetosonic speed from the MAS MHD model.
@@
For the notable 2017 September 10 global event, the transmission of the EUV wavefront by two polar coronal holes has been reported in previous researches (e.g., Liu et al., 2018; Hu et al., 2019). Before entering the south polar coronal hole, the wavefront speed increases with the altitude from 600 km s-1 to 1000 km s-1 (see Figure 6(b)–(f)). From the inbound boundary of the coronal hole (near position angle 210°), only the wavefront at altitudes of 1.05–1.15 is visible. Inside the coronal hole, the apparent wavefront speed is very high, and ranges from 1000 km s-1 to even 3000 km s-1, which is comparable to the estimate of Liu et al. (2018). After leaving the south polar coronal hole, the wavefront undergoes a gradual deceleration before it collides with the wavefront transmitted by the north polar coronal hole (see Hu et al., 2019). Behind the outbound boundary of the south polar coronal hole (near position angle 140°), the apparent wavefront speed is still km s-1 and the Mach number can be nearly 3 (see Figure 6(b)–(d)). Considering the inclination angle of ° for the transmitted wavefront, the actual speed can be below 500 km s-1 and the Mach number can be below 2 (see the cyan curves in Figure 6(b)–(d)). Inside the coronal hole the inclination angle is not estimated because the wavefront is too diffuse to obtain a reliable angle with visible inspection. If we assume the inclination angle is the same as the transmitted wavefront behind the coronal hole, the actual wavefront speed inside the coronal hole can also be lower than the apparent speed. In the initial stage, the inclination angle of the primary wavefront is °, which means the wavefront is less inclined than those of other events and the actual speed is approximately equal to the apparent speed.
@@ 3.3 Coupling of the CME Boundary and the EUV Wave
For the 2011 September 22 and 2013 November 7 events, we can see the angular distance between the EUV wavefront and the CME boundary is close to zero (see Column (6) of Table 2), which suggests that the two features are coupled until the EUV wave fades out. For the other four events, the angular distance between the EUV wavefront and the CME boundary is significantly greater than 0°, which will be discussed separately below.
@@
Due to the large cadence of the SOHO/LASCO C2 images, it is not possible to track the CME boundary during the whole propagation of the EUV wavefront. For the 2011 September 22 and 2013 November 7 events, the angular distance between the EUV wavefront and the CME boundary is approximately zero, when the wavefront fades out at all altitudes. Although the white-light CME structure below 2 cannot be observed by SOHO/LASCO C2, the CME boundary is almost radial as illustrated in Figure 9 and can extend straight downward to below 2 . This is consistent with that the legs or the overlying loops of the CME boundary are stretched to straight during the rising and expansion of a CME structure. It is indicated that the EUV wavefront is still coupled with the CME boundary when the wavefront vanishes. This suggests during the whole propagation, the EUV wavefront is not propagating freely but is continuously driven by the expanding CME in the two events.